Circular No. 9051

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 221P/2009 L1 (LINEAR) L. Elenin, Lyubertsy (Moscow region), Russia, reports the recovery of

comet P/2002 JN_{16} (cf. IAUC 7907) on CCD images obtained remotely with a 0.36-m f/3.8 Maksutov-Newtonian telescope near Mayhill, NM, USA (first observation tabulated below). The recovery was confirmed by M. Schwartz with a 0.81-m f/7 reflector near Rio Rico, AZ, USA, whose images clearly show a tail of length ∼ 35′′ in p.a. 248^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
June 1.39819 23^{h}26^{m}08.^{s}23 +8^{o}37′27′′.6 20.5 Elenin

The correction to the predictions on MPC 56802 and in the 2008/2009 Comet Handbook is ΔT = −0.2 day. Additional information, the available astrometry, the following revised orbital elements, residuals, and an ephemeris appear on CBET 1831. The permanent number 221P has been assigned to this comet (cf. MPC 66145).

Epoch = 2009 Jan. 9.0 TT

T = 2009 Jan. 24.88878 TT  ω = 39.^{o}69673
e = 0.4872673  Ω = 230.03375 }2000.0
q = 1.7836977 AU  i = 11.41862

a = 3.4788059 AU n^{o} = 0.15190035 P = 6.49 years

COMETS C/2008 E10, C/2008 Y16, AND C/2008 Y17 (STEREO) Further to IAUCs 9025 and 9036, additional slightly diffuse and some-

what elongated Kreutz sungrazers have been found on STEREO/SECCHI HI-1A images (the elongation is partly due to the long exposure times, ∼ 40 min). Approximate peak magnitudes: C/2008 E10, 11–12; C/2008 Y16, 10–11; C/2008 Y17, 10. K. Battams notes that Kreutz comets in the HI-1 images generally initially appear out of the background noise at mag ∼ 13.5, and they are still in the process of brightening as they leave the fieldof-view; he adds that their rates of initial brightening vary — some making very abrupt appearances, while others gradually come into view.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 E10 Mar. 7.850 1^{h}01.^{m}8 + 2^{o}32′ HI* AW 2009-F31
C/2008 Y16 Dec. 31.794 22 05.1 −15 04 HI* AW 2009-F31
C/2008 Y17 Dec. 31.905 22 07.4 −15 32 HI* KB 2009-F31
2009 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green